# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ III-C The score vectors and the related FIM in $\mathcal{P}_{\bm{\eta},g}$
$\begin{split}E\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{Q}\}&=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})\mathrm{vec}(E\{\mathbf{u}\mathbf{u}^{T}|\mathcal{Q}\})-\mathrm{vec}(\mathbf{I}_{m})\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathbf{I}_{m})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\varphi_{0}(\mathcal{Q})-m\right)\\ &=E\{\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})|\mathcal{Q}\}.\end{split}$ (83)
Consequently, the condition in (79) follows. Hence, the condition of Lemma II.2, equivalent to (77) in this setting, is verified.
To prove the second part of Proposition IV.1, we need to show that, if $S(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$, then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. From (82), we have that:
or equivalently $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$ iff $\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\in\text{Ker}\;\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})=\{a\cdot\mathrm{vec}(\mathbf{D}_{S}^{\mathbf{V}_{S,0}}),a\in\mathbb{R}\}$ where the matrix $\mathbf{D}_{S}^{\mathbf{V}_{S,0}}$ is formally defined in (157) of .6. As already noted in [10, Th. 3.1], this is a direct consequence of the Property P4 of the matrix $\mathbf{M}_{S}^{\mathbf{V}_{S,0}}$ (see appendix .6). In fact, for $S\equiv S_{d}$, we have that $\mathbf{D}_{S_{d}}^{\bm{\Sigma}_{0}}=m^{-1}|\bm{\Sigma}_{0}|^{1/m}\bm{\Sigma}_{0}^{-1}=m^{-1}\mathbf{V}_{S,0}^{-1}$. Consequently, we immediately have that $\mathrm{vec}(\mathbf{V}_{S_{d},0}^{-1})\in\text{Ker}\;\mathbf{M}_{S_{d}}^{\mathbf{V}_{S,0}}\cap\mathrm{vec}(\mathcal{S}_{m}^{\mathbb{R}})$ and then $\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})=\mathbf{0}$. ∎
@@
As a side result of this proof and of the fact that $\mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=\mathbf{0}$ and $\mathbf{I}_{\bm{\mu}_{0},s_{0}}=\mathbf{0}$, it is immediate to verify that knowing or not knowing the location vector $\bm{\mu}_{0}$ has no impact on the efficiency losses when we do parametric or semiparametric inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$. Moreover, since $\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{g_{0}})=\mathbf{0}$, the lack of knowledge of the density generator does not induce any loss of efficiency w.r.t. $\bm{\mu}_{0}$ (see e.g. the discussion in [22]).
Roughly speaking Proposition IV.1 tells us that:
@@
  • When performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the scale $s_{0}$ (while knowing the density generator $g_{0}$) or not knowing neither the scale $s_{0}$ nor the density generator $g_{0}$, leads to the same efficiency losses.
@@
  • If the scale $S_{d}(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$ is adopted, when performing inference on $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S_{d}}^{\mathbb{R}}$, knowing or not knowing the scale $s_{0}$ and/or the density generator $g_{0}$, does not lead to any asymptotic efficiency loss w.r.t. the case in which $s_{0}$ and/or $g_{0}$ are perfectly known. For this reason, the scale $S_{d}(\bm{\Sigma}_{0})$ is called ”canonical” in [10].
@@
Let us now focus our attention on the parametric model $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ in (48) in which we have two finite-dimensional nuisance parameters: the scale $s_{0}$ and the parameters of the density generator $\bm{\zeta}_{0}\in\Delta$. The difference between $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ and the semiparametric model $\mathcal{P}_{\bm{\eta},g}$ in (40) is in the fact that in $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ the functional form of the density generator $g$ is assumed to be known up to a finite-dimensional vector of parameters while $g$ is fully unknown in $\mathcal{P}_{\bm{\eta},g}$.
Proposition IV.2.
The efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\eta},g}$ in (40) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ in (48).
@@
Let us start by proving that the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}_{\gamma,s,\bm{\zeta}}$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$ in (46). To this end, we just need to apply the Lemma II.2 by substituting the infinite-dimensional nuisance tangent space $\mathcal{T}_{3}$ with the finite-dimensional one generated by the score vector $\mathbf{s}_{\bm{\zeta}_{0}}$ in (11), i.e.
$\mathcal{H}\supseteq\mathcal{T}_{\bm{\zeta}_{0}}=\mathrm{Span}\{[\mathbf{s}_{\bm{\zeta}_{0}}]_{1},\ldots,[\mathbf{s}_{\bm{\zeta}_{0}}]_{p}\}.$ (85)
Consequently, we need to show that:
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\triangleq\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\perp\mathcal{T}_{\bm{\zeta}_{0}},$ (86)
or, equivalently, that:
$\Pi(\bar{\mathbf{t}}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\zeta}_{0}})=E_{0}\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}\mathbf{I}_{\bm{\zeta}_{0}}^{-1}\mathbf{s}_{\bm{\zeta}_{0}}=\mathbf{0},$ (87)
where $\mathbf{I}_{\bm{\zeta}_{0}}\triangleq E_{0}\{\mathbf{s}_{\bm{\zeta}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}$ and $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$, from (82), can be expressed as:
$\begin{split}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}&=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ \mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}-\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}})\end{array}\right)\\ &=\left(\begin{array}[]{c}\mathbf{s}_{\bm{\mu}_{0}}\\ 2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathcal{Q}\varphi_{0}(\mathcal{Q})[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathrm{vec}(\mathbf{u}\mathbf{u}^{T})-m^{-1}\mathrm{vec}(\mathbf{I}_{m})\right)\end{array}\right).\end{split}$ (88)
We can also introduce the efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ (c.t. (23)) in the parametric model $\mathcal{P}_{\bm{\gamma},s}$ in (46) (c.t. $\mathcal{P}_{1,2}$ in (8)) for $\bm{\gamma}_{0}$ in the presence of the finite-dimensional nuisance terms $s_{0}$ as:
$\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})\triangleq E_{0}\left\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\bar{\mathbf{t}}_{\bm{\gamma}_{0}}^{T}\right\}=\left(\begin{array}[]{cc}\mathbf{I}_{\bm{\mu}_{0}}&\mathbf{0}\\ \mathbf{0}&\bar{\mathbf{I}}(\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})|s_{0})\end{array}\right),$ (89)
whose block-diagonal structure follows once again from the properties of $\mathbf{u}\sim\mathcal{U}(S_{\mathbb{R}}^{m-1})$ and from the independence of $\mathbf{u}$ with $\mathcal{Q}$.
Clearly, the equality in (87) is verified iff $E_{0}\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}=\mathbf{0}$. This condition can be shown to hold using the mutual independence of $\mathcal{Q}$ and $\mathbf{u}$ and from the properties of $\mathbf{u}$. In fact, from the expression of $\bar{\mathbf{t}}_{\bm{\gamma}_{0}}$ given in (88), we have that:
$\begin{split}E_{0}&\{\bar{\mathbf{t}}_{\bm{\gamma}_{0}}\mathbf{s}_{\bm{\zeta}_{0}}^{T}\}\\ &=\left(\begin{array}[]{c}\bm{\Sigma}^{-1/2}_{0}E\left\{\frac{\sqrt{\mathcal{Q}}\varphi_{0}(\mathcal{Q})}{g_{\bm{\zeta}_{0}}(\mathcal{Q})}\bm{\nabla}^{T}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q})\right\}E\{\mathbf{u}\}\\ 2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]E\left\{\frac{\mathcal{Q}\varphi_{0}(\mathcal{Q})}{g_{\bm{\zeta}_{0}}^{2}(\mathcal{Q})}\bm{\nabla}_{\bm{\zeta}}g_{\bm{\zeta}_{0}}(\mathcal{Q})\bm{\nabla}_{\bm{\zeta}}^{T}g_{\bm{\zeta}_{0}}(\mathcal{Q})\right\}\left(\mathrm{vec}(E\{\mathbf{u}\mathbf{u}^{T}\})-m^{-1}\mathrm{vec}(\mathbf{I}_{m})\right)\end{array}\right)=\mathbf{0}.\end{split}$ (90)
since $E\{\mathbf{u}\}=\mathbf{0}$ and $E\{\mathbf{u}\mathbf{u}^{T}\}=m^{-1}\mathbf{I}_{m}$. Finally, the statement of Proposition IV.2 follow immediately from the fact that $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$, as shown in Proposition IV.1. ∎
@@
Roughly speaking Proposition IV.2 tells us that, if the scale $s_{0}$ is an unknown parameter, when performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the parameters $\bm{\zeta}_{0}$ of the density generator $g_{\bm{\zeta}_{0}}$ or not knowing the whole functional form $g$, leads to the same efficiency losses.
@@
Propositions IV.1 and IV.2 give us a clear picture of the efficiency relationships among the various sub-models of elliptical distributions in terms of Fisher information matrices. It would now be interesting to obtain explicit closed-form expressions for the inverses of these FIM matrices in order to obtain the related information bounds. This will be the objective of the next section.
V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\eta},g}$
@@
The purpose of this section is to provide the counterpart of the Propositions IV.1 and IV.2 in terms of information bounds.
@@
A clarification is in order before continuing. In the theory of parametric estimation, the best known information bound is the Cramér-Rao Bound (CRB) which coincides with inverse of the related FIM as shown e.g. [17, Chap. 5, Sects 5 and 6]. Consequently, for the different parametric submodels of the canonical semiparametric model in $\mathcal{P}_{\bm{\eta},g}$ in (40), we can derive the related information/CR bounds as:
@@
  • CRB for the parameters of interest $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ in the presence of the finite-dimensional nuisance $s_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\gamma},s}$ in (46):
    $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})^{-1},$ (91)
    where $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ is the efficient FIM as defined in (89).
@@
  • CRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of two finite-dimensional nuisance terms, $s_{0}$ and $\bm{\zeta}_{0}$ in the parametric submodel $\mathcal{P}_{\bm{\gamma},s,\bm{\zeta}}$ in (48):
@@
    as shown in Proposition IV.2.
When we move to the semiparametric case, the CRB can no longer be defined as in the classical parametric case. Remarkably, the Hájek-Le Cam convolution theorem (see e.g. [1, Sect. 3.3, Theo. 2]) provides the right theoretical framework to unify the concept of information bound in the parametric, semiparametric and non-parametric case. A formal presentation of this theorem would lead us too far from the main purpose of this article. Therefore, below we will simply define the “Semiparametric CRB (SCRB)” as the information bound obtained as the inverse of the efficient semiparametric FIM (SFIM) (for an in-depth discussion about this point, we refer the readers to [1, Chap. 3]). Specifically, for the semiparametric canonical model $\mathcal{P}_{\bm{\eta},g}$ in (40), we have that the SCRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of a finite-dimensional $s_{0}$ and of in infinite-dimensional $g_{0}$ nuisance terms is given by:
$\mathrm{SCRB}(\bm{\gamma}_{0}|s_{0},g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})^{-1},$ (93)
where $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},g_{0})$ is the efficient Semiparametric FIM as defined in (22).
The following two subsections are organized as follows: in subsect. V-A, an explicit expression for the parametric CRB of $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ in the presence of the nuisance parameter $s_{0}$ is provided, while the subsect. V-B proposes the counterpart of the Propositions IV.1 and IV.2 in terms of parametric and semiparametric information bounds.
V-A Explicit expression for $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}_{\bm{\gamma},s}$
To calculate (in a closed-form not containing matrix inverses) the $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in (91), we will exploit the expression of the CRB on $\bm{\nu}_{0}\triangleq(\bm{\mu}_{0}^{T},\mathrm{vecs}(\bm{\Sigma})_{0}^{T})^{T}$ deduced from the inverse of the FIM $\mathbf{I}_{\bm{\nu}_{0}}$. To this end, let us consider the inverse, say $\mathbf{w}^{-1}$, of the diffeomorphism $\mathbf{w}$ given in (50). It can be explicitly obtained as:
@@
The parametric CRB for $\bm{\gamma}_{0}$ in $\mathcal{P}_{\bm{\gamma},s}$ is given by
@@
$\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\frac{s_{0}}{\beta(g_{0})}\mathbf{V}_{S,0}$ (97)

### v3

@@ III-A Essentials on RES distributions
$\begin{split}\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}^{u}_{g_{0}})&=E\{\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{Q}\}=_{d}2^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\left(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\mathrm{vec}(E\{\mathbf{u}\mathbf{u}^{T}|\mathcal{Q}\})-\mathrm{vec}(\mathbf{I}_{m})\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathbf{I}_{m})\left(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-m\right)\\ &=(2m)^{-1}\mathbf{M}_{S}^{\mathbf{V}_{S,0}}\mathrm{vec}(\mathbf{V}_{S,0}^{-1})\left(\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-m\right)=\Pi(\mathbf{s}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}|\mathcal{T}_{s_{0}}),\end{split}$ (88)
where we recall that $\mathcal{T}_{s_{0}}\subset\mathcal{T}^{u}_{g_{0}}$. As a consequence of the previous outcomes and of the relation in (82), we have that the efficient semiparametric score vector $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$ can be expressed in three equivalent forms:
$\begin{split}\bar{\mathbf{s}}_{\bm{\gamma}_{0}}&\triangleq\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}_{s_{0}}\\ &=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}^{u}_{g_{0}})\perp\mathcal{T}^{u}_{g_{0}},\\ &=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}}+\mathcal{T}^{c}_{\bar{g}_{0}}),\end{split}$ (89)
and then $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|g_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$.
We note in passing that the equality $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$ can be obtained by exploiting Lemma II.2. Specifically, in this context, Condition (28) of Lemma II.2 can be expressed as:
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})]\perp\mathcal{T}^{c}_{\bar{g}_{0}},$ (90)
that follows immediately from (89) and from the fact that $\mathcal{T}^{c}_{\bar{g}_{0}}\subset\mathcal{T}^{u}_{g_{0}}$.
To conclude, the expression of $\bar{\mathbf{I}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ can be derived from the efficient score $\bar{\mathbf{s}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}$ in (61) and (88) as:
$\bar{\mathbf{s}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=_{d}2^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\mathbf{M}_{S}^{\mathbf{V}_{S,0}}[\mathbf{V}_{S,0}^{-1/2}\otimes\mathbf{V}_{S,0}^{-1/2}]\mathrm{vec}(\mathbf{u}\mathbf{u}^{T}-m^{-1}\mathbf{I}_{m}),$ (91)
from which (84) of $\bar{\mathbf{I}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\triangleq E\{\bar{\mathbf{s}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}\bar{\mathbf{s}}_{\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}^{T}\}$ follows directly. [^6] The calculation is based on the independence between $\mathcal{Q}$ and $\mathbf{u}$ along with the identities $E\{{\rm vec}(\mathbf{u}\mathbf{u}^{T}){\rm vec}^{T}(\mathbf{u}\mathbf{u}^{T})\}=\frac{1}{m(m+2)}\left({\bf I}_{m^{2}}+{\bf K}_{m}+{\rm vec}(\mathbf{I}_{m}){\rm vec}^{T}(\mathbf{I}_{m})\right)$ and $\mathbf{K}_{m}\mathbf{D}_{m}=\mathbf{D}_{m}$ [25, p. 49] in (55) We note, in passing that similar expressions have been derived in [11, eqs. (8) and (9)], by using the already-recalled asymptotic representation of the projection operator. ∎
@@
As a side result of this proof and of the fact that $\mathbf{I}_{\bm{\mu}_{0},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})}=\mathbf{0}$ and $\mathbf{I}_{\bm{\mu}_{0},s_{0}}=\mathbf{0}$, it is immediate to verify that knowing or not knowing the location vector $\bm{\mu}_{0}$ has no impact on the efficiency losses when we do parametric or semiparametric inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$. Moreover, since $\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{g_{0}}^{u})=\Pi(\mathbf{s}_{\bm{\mu}_{0}}|\mathcal{T}_{g_{0}}^{c})=\mathbf{0}$, the lack of knowledge of the density generator does not induce any loss of efficiency w.r.t. $\bm{\mu}_{0}$ (see e.g. the discussion in [24]).
@@
Roughly speaking, Proposition IV.1 tells us that when performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the scale $s_{0}$ (while knowing the constrained density generator $\bar{g}_{0}$) or not knowing the density generator at all $g_{0}$, leads to the same efficiency losses.
@@
Remark 1: In [11], it has been show that, if the scale function $S_{d}(\bm{\Sigma}_{0})\triangleq|\bm{\Sigma}_{0}|^{1/m}$ is adopted, then the parametric FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ in the parametric model $\mathcal{P}_{\bm{\gamma}}$ in (43) is equal to the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$ for $\mathcal{P}_{\bm{\eta},\bar{g}}$. In other worlds, if the scale $S_{d}(\bm{\Sigma}_{0})=|\bm{\Sigma}_{0}|^{1/m}$ is adopted, when performing inference on $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S_{d}}^{\mathbb{R}}$, knowing or not knowing the scale $s_{0}$ and/or the density generator $g_{0}$, does not lead to any asymptotic efficiency loss w.r.t. the case in which $s_{0}$ and/or $g_{0}$ are perfectly known. For this reason, the scale $S_{d}(\bm{\Sigma}_{0})$ is called “canonical” in [11]. A sketch of the proof of this result is recalled in Appendix .9.
@@
Let us now focus our attention on the parametric model $\mathcal{P}^{c}_{\gamma,s,\bm{\zeta}}$, whose sub-model is given in (47), in which the finite-dimensional nuisance vector is represented by the parameters $\bm{\zeta}_{0}\in\Delta$ of the unconstrained density generator.
Corollary IV.1.
The efficient parametric FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ in (45) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}^{c}_{\gamma,s,\bm{\zeta}}$ in (47).
@@
In Proposition IV.1, we showed that $\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}^{c}_{\bar{g}_{0}}$. Moreover since, as discussed in Appendix .8, $\mathcal{T}^{c}_{\bm{\zeta}_{0}}\subset\mathcal{T}^{c}_{\bar{g}_{0}}$, we trivially have that $\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{s_{0}})\perp\mathcal{T}^{c}_{\bm{\zeta}_{0}}$. Consequently, the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bm{\zeta}_{0})$ in $\mathcal{P}^{c}_{\gamma,s,\bm{\zeta}}$ is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ in (45). ∎
@@
Roughly speaking, Corollary IV.1 tells us that, if the scale $s_{0}$ is an unknown parameter, when performing inference on the shape matrix $\mathbf{V}_{S,0}\in\mathcal{S}_{m,S}^{\mathbb{R}}$, not knowing the parameters $\bm{\zeta}_{0}$ of the density generator $\bar{g}_{\bm{\zeta}_{0}}$ or not knowing the whole functional form $\bar{g}$, leads to the same efficiency losses.
@@
Proposition IV.1 and Corollary IV.1 give us a clear picture of the efficiency relationships among the various sub-models of elliptical distributions in terms of Fisher information matrices. It would now be interesting to obtain explicit closed-form expressions for the inverses of these FIM matrices in order to obtain the related information bounds. This will be the objective of the next section.
V Parametric and semiparametric information bounds in $\mathcal{P}_{\bm{\gamma},g}$ and $\mathcal{P}_{\bm{\eta},\bar{g}}$
@@
The purpose of this section is to provide the counterpart of the Propositions IV.1 in terms of information bounds.
@@
A clarification is in order before continuing. In the theory of parametric estimation, the best known information bound is the Cramér-Rao Bound (CRB) which coincides with inverse of the related FIM as shown e.g. [19, Chap. 5, Sects 5 and 6]. Consequently, for the different parametric sub-models of the semiparametric model in $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40), we can derive the related information/CR bounds as:
@@
  • CRB for the parameters of interest $\bm{\gamma}_{0}=(\bm{\mu}_{0}^{T},\underline{\mathrm{vecs}}(\mathbf{V}_{S,0})^{T})^{T}$ in the presence of the finite-dimensional nuisance $s_{0}$ in the parametric sub-model $\mathcal{P}^{c}_{\bm{\gamma},s}$ in (45):
    $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0})^{-1}.$ (92)
@@
  • CRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of two finite-dimensional nuisance terms, $s_{0}$ and $\bm{\zeta}_{0}$ in the parametric sub-model $\mathcal{P}^{c}_{\bm{\gamma},s,\bm{\zeta}}$ in (47):
@@
    as shown in Corollary IV.1.
When we move to the semiparametric case, the CRB can no longer be defined as in the classical parametric case. Remarkably, the Hájek-Le Cam convolution theorem (see e.g. [1, Sect. 3.3, Theo. 2]) provides the right theoretical framework to unify the concept of information bound in the parametric, semiparametric and non-parametric case. A formal presentation of this theorem would lead us too far from the main purpose of this article. Therefore, below we will simply define the “Semiparametric CRB (SCRB)” as the information bound obtained as the inverse of the efficient semiparametric FIM (SFIM) (for an in-depth discussion about this point, we refer the readers to [1, Chap. 3]). Specifically, for the semiparametric model $\mathcal{P}_{\bm{\eta},\bar{g}}$ in (40), we have that the SCRB for the parameters of interest $\bm{\gamma}_{0}$ in the presence of a finite-dimensional $s_{0}$ and of an infinite-dimensional $\bar{g}_{0}\in\overline{\mathcal{G}}$ nuisance terms is given by:
$\mathrm{SCRB}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})=\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})^{-1},$ (94)
where $\bar{\mathbf{I}}(\bm{\gamma}_{0}|s_{0},\bar{g}_{0})$ is the efficient Semiparametric FIM as defined in (22).
The following two subsections are organized as follows: in subsect. V-A, an explicit expression for the parametric CRB of $\bm{\gamma}_{0}$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ in the presence of the nuisance parameter $s_{0}$ is provided, while the subsect. V-B proposes the counterpart of the Proposition IV.1 in terms of parametric and semiparametric information bounds.
V-A Explicit expression for $\mathrm{CRB}(\bm{\gamma}_{0}|s_{0})$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$
Let us consider the inverse, say $\mathbf{w}^{-1}$, of the diffeomorphism $\mathbf{w}$ given in (50). It can be explicitly obtained as:
@@
The parametric CRB for $\bm{\gamma}_{0}$ in $\mathcal{P}^{c}_{\bm{\gamma},s}$ is given by
@@
$\mathrm{CRB}(\bm{\mu}_{0})\triangleq\mathbf{I}_{\bm{\mu}_{0}}^{-1}=\frac{s_{0}}{\beta(\bar{g}_{0})}\mathbf{V}_{S,0}$ (98)